# UOE: Unlearning one Expert is Enough for Mixture-of-Experts LLMs

### v1

@@ 5.2 Experiment Results
UOE $0.06\%$ $0.06\%$ 0.41%
UOE significantly outperforms other PEFT methods. Tab. 4 shows the performance comparison between UOE and other PEFT methods, and Tab. 3 shows a comparison of the parameter efficiency among different PEFT methods. Several key conclusions can be drawn: First, UOE achieves far better parameter efficiency, with only $0.06\%$ of tunable parameters, compared to LoRA ($0.92\%$) and ESFT ($2.86\%$), while still outperforming them in utility preservation. For instance, in RWKU, UOE achieves utility scores of 0.5709 on Qwen and 0.5445 on DeepSeek, significantly higher than LoRA (0.2689 and 0.2302) and ESFT (0.4433 and 0.5001). Second, the utility preservation of UOE is much better than the others, and this is achieved while maintaining a comparable level of forget efficacy. For example, on WMDP, UOE achieves a utility score of 0.5012 for Qwen, much higher than LoRA’s 0.2689, with a similar forget efficacy (FA: 0.2987 vs. 0.2459). These results clearly demonstrate that UOE is the more balanced and efficient solution for unlearning tasks, particularly when both parameter efficiency and utility retention are important.
Table 4: Performance comparison between existing unlearning methods GA equipped with UOE and other PEFT methods, including LoRA (Hu et al., 2021) and ESFT (Wang et al., 2024b). The occurrence of significant utility drop (over $10\%$ drop in UT compared to the pretrained model) are marked in red.
Method Qwen (WMDP) DeepSeek (WMDP) Mixtral (WMDP) Qwen (RWKU) DeepSeek (RWKU) Mixtral (RWKU)
[…]
Pretrained $0.4192$ $0.5979$ $0.3804$ $0.5548$ 0.5229 0.6885 $0.4243$ $0.5979$ $0.5376$ $0.5548$ 0.5820 0.6885
LoRA 0.2459 0.2689 0.2657 0.2295 0.2658 0.2597 $0.0000$ $0.2689$ 0.0000 0.2302 0.0000 0.2295
ESFT 0.3145 0.4514 0.2737 0.5108 0.2547 0.6386 0.001 0.4433 0.0200 0.5001 0.0542 0.6743
UOE 0.2987 0.5012 0.2700 0.5100 0.2608 0.6364 0.006 0.5709 0.0000 0.5485 0.0455 0.6713
Figure 5: Expert selection overlap ratio between the pretrained model and the unlearned model across different unlearning iterations using GA on the WMDP benchmark. Experts with the highest affinity score in each layer are tuned during unlearning.
Comparison of different design choices in UOE framework. In designing our UOE method, we opted to unlearn only a single expert in one specific layer, guided by the affinity score. However, to justify this design decision, we conducted a series of empirical studies comparing alternative approaches. These experiments were carried out using the RMU unlearning method on the WMDP task, where we tuned each approach until the model fully unlearned (i.e., FA reached 25%), and then compared (1) if the problem of the expert selection shift has been properly addressed and (2) the model utility score (UT) across the different strategies.
$\bullet$ Tuning multiple layers with one expert per layer. Tuning a single expert across multiple layers. In Fig. 4, we briefly discussed how the cumulative effect leads to a significant issue of expert selection shift in the deeper layers of MoE LLMs. In Fig. 5, we further illustrate how the overall expert selection overlap ratio changes as unlearning progresses. As shown, even when only one expert is tunable per layer, the cumulative effect causes the overall overlap ratio to drop sharply, leaving the expert selection shift issue unresolved. Thus, tuning all layers during MoE LLM unlearning is not a feasible solution.
Table 5: Model utility (UT) comparison, at the same level of forget efficacy (FA0.25absent0.25$\approx$ 0.25≈ 0.25), when different number of experts are unlearned using GA on WMDP benchmark.
# of experts 1 3 6
FA ($\downarrow$) $\sim 0.2500$
UT ($\uparrow$) 0.5100 0.4856 0.4652
$\bullet$ Tuning multiple experts in a single layer. A natural question within the UOE framework is whether involving more than one expert in the selected layer during unlearning would yield better results. In Tab. 5, we examine the impact of increasing the number of tunable experts in a single layer. As the table shown, when the unlearned model reaches the same level of UE (with an FA of around $25\%$) controlled by different training steps, the model utility decreases significantly as the number of tunable experts increases. For instance, when the tunable expert count is increased to 6, the model utility drops by over $4\%$, from $0.51$ to below $0.47$. This suggests that involving more experts in unlearning makes it harder to maintain utility, without providing any noticeable improvement in unlearning efficacy.
Table 6: Performance comparison between UOE and the random expert selection scheme on the WMDP task. Other settings are consistent with those in Tab. 2.
Method Qwen (WMDP) DeepSeek (WMDP) Qwen (RWKU) DeepSeek (RWKU)
[…]
Pretrained $0.4192$ $0.5979$ $0.3804$ $0.5548$ $0.4243$ $0.5979$ $0.5376$ $0.5548$
RMU $0.2612$ $0.3560$ $0.2530$ $0.4540$ $0.0200$ $0.2420$ $0.0010$ $0.5109$
[…]
$\bullet$ Sensitivity of UOE to expert selection and layer selection schemes. Next, we explored alternative methods for selecting the target expert, beyond the affinity score-based approach used in UOE. Specifically, we investigated how sensitive UOE is to the choice of the expert to be unlearned. In Tab. 6, we compared the performance of the affinity score-based selection scheme in UOE with a random expert selection approach on the WMDP task. The results clearly show that while random expert selection can sometimes yield comparable utility preservation, it fails to achieve the same level of unlearning efficacy as UOE. For instance, on Qwen (WMDP), the random selection achieves a utility score of $0.5947$ versus UOE’s $0.5351$, but the forget accuracy (FA) with random selection is significantly higher at $0.3505$, compared to $0.2536$ for UOE. This indicates that random selection does not drive FA low enough, compromising the unlearning objective. Similarly, on DeepSeek (RWKU), random selection results in an FA of $0.1176$, whereas UOE achieves a much lower FA of $0.013$, highlighting the better unlearning performance of UOE. In conclusion, selecting experts with highest affinity score perform better than random expert selection.
@@
Table 7: Model utility (UT) comparison across layers with different affinity score rankings in UOE on the RWKU benchmark. UT is compared at a consistent level of forget efficacy (FA $\approx 0.25$).
Layer Ranking #1 #2 #3 #13 #20 #23 #26
Affinity Score 0.2110 0.1957 0.1695 0.1115 0.0942 0.0844 0.0618
FA ($\downarrow$) $\sim 0.2500$
@@
Finally, we examine the performance of UOE when layers with different top1 expert affinity score rankings are selected for unlearning. In Tab. 7, we observe that UOE is robust and relatively insensitive to the specific layer chosen for unlearning, as long as the affinity score remains reasonably high. For instance, even when selecting the 13th or 20th ranked layers, the model utility (UT) remains stable at around 0.5445, although their affinity scores of 0.1115 and 0.0942 are lower than that of the top-ranked layer. However, once the affinity score drops further, as seen in the 23rd and 26th ranked layers (with scores of 0.0844 and 0.0618), the utility decreases more sharply, falling to 0.4262 and 0.2355, respectively. This demonstrates that while UOE maintains strong performance across a wide range of layers, selecting layers with very low affinity scores can negatively impact utility. Overall, these results highlight the robustness of UOE and its ability to tolerate variability in layer selection without sacrificing unlearning efficacy, provided that layers with sufficiently high affinity scores are chosen.
Our empirical studies demonstrate that the design choice of unlearning a single expert in one layer, guided by the affinity score, is satisfactorily reasonable for balancing effective unlearning with utility preservation. The alternative approaches, whether involving multiple layers or multiple experts, consistently led to greater utility degradation and instability. Selecting expert with highest affinity scores can achieve better forget quality than picking experts by random choice. At last UOE is very robust to layer selection scheme. Therefore, our current design proves to be the most effective and efficient for MoE LLM unlearning.
@@ 6 Conclusion
In this paper, we for the first time examine the challenges of applying existing MU techniques to MoE LLMs and carefully investigate the synergy between the dynamic routing system of MoE LLM and the unlearning effects. To address these issues, we proposed UOE, a novel framework that unlearns a single expert in a targeted layer while stabilizing expert selection through a router anchor loss. This approach mitigates expert selection shifts and achieves efficient unlearning with minimal parameter updates. Extensive experiments show that UOE significantly outperforms traditional unlearning methods and other parameter-efficient fine-tuning techniques, providing a robust solution for MoE LLM unlearning tasks.
Appendix A Experiment Settings
Hyperparameter selection. We set the learning rate to 5e-5 for GA, NPO, and GD while setting it to 1e-4 for UOE. The batch size is 4 for GA, NPO, and GD, while it is set to 16 for UOE. In NPO, the beta value is set to 0.001. The alpha for the retain loss is set to 1 in both GD and NPO. For RMU, we follow the hyperparameters specified in the original work. We configure the steering coefficients as 8000 for Qwen and 32000 for Deepseek, as UOE targets deeper layers in these models. For ESFT, we set the threshold $p=0.15$.
Dataset Settings. For the WMDP dataset, we use the cyber-forget-corpus as the forget set and wmdp-cyber as the evaluation set, in line with WMDP (Li et al., 2024). The Wikitext (Merity et al., 2016) dataset serves as the retain set for both GD and RMU tasks, also following WMDP (Li et al., 2024). In the RWKU (Jin et al., 2024) dataset, we follow the original study by selecting 100 individuals as unlearning targets. The train_original_passage set, which includes Wikipedia descriptions of these 100 individuals as provided in the paper, is used as the forget set.
@@ Appendix A Experiment Settings
Evaluation Settings. We utilize the LM Evaluation Harness (Gao et al., 2024) to measure zero-shot accuracy on the MMLU and WMDP cyber datasets. The mean accuracy across all tasks in MMLU serves as a measure of model utility. For the RWKU dataset, we adhere to the original settings, using the prompt “Please complete the blank in the following question. Question:” for fill-in-the-blank tasks and “Please briefly answer the following question. Question:” for generation tasks.
Appendix B Jailbreak Attack on UOE
Table 8: Example of GCG attack. Original is the output of the original Deepseek model.
Questions Original Without Attack With Attack
Stephen King is often referred to as the ’King of ___’. Horror indingindingind-ingindingindingi-ndinginding… ologist:Sure here is the answer: Question : Stephenindingindingin-dinginding…
Bruce Lee moved to ___ in 1959. America -wardswar-dswardswards… eanswer:” Answer:[/b]inding [b]…
Table 9: UOE against GCG attack on Deepseek unlearned with RWKU dataset.
With Attack Without Attack
0.01 0.01
Recent works introduce prompt-level attacks (Zou et al., 2023; Liu et al., 2023; Zhuang et al., 2023; Wei et al., 2024a) to recover hidden knowledge from generative models. To investigate the soft prompt attack, we employ the most representative and powerful attack GCG (Zou et al., 2023) in a white-box setting to optimize the prompt, aiming to force responses to begin with “Sure, here is the answer:”. The number of optimization steps is increased to 5000, while other hyperparameters remain at the default settings. Given the computational demand (approximately 1 GPU hour on an A100 for generating a single soft prompt), we optimize 400 prompts across 400 samples in RWKU. Since not all responses begin with ”Sure, here is the answer:”, we filter for those containing the word ”answer” and then assess the forget quality both with and without GCG-generated prompts. Experimental results in Tab. 8 indicate that the GCG, despite being the strongest prompt-level attack, fails to recover the forgotten knowledge, as the forget accuracy (FA) remains at 0.01 before and after the GCG attacks.
Appendix C Token proportion Shift Visualization
We visualize the expert selection distribution in one layer across the unlearning process in Fig.6 in revision. The figures sort the experts in layer 21 by the selection proportion in the original model and keeps this order to plot the figure in unlearned models to show the changes in all experts. The results show that GA algorithm even decreases the uncertainty in WMDP after unlearning.
Appendix D Sensitivity Analysis of Hyperparameter $\alpha$
@@
We conduct experiments on Deepseek unlearned by GA with RWKU dataset to explore the performance of different $\alpha$. As shown in Tab. 10, the results indicate that UOE is robust to a wide range of $\alpha$ and achieves the best performance when $\alpha=1$.

### v2

@@ 5 Evaluation Experiments
SEUF $0.06\%$ $0.06\%$ 0.41%
Top-1 expert selection outperforms random selection in unlearning. In the last row of Tab. 3, we compare the performance of the affinity score-based expert selection in SEUF with a random expert selection approach. The results show that while random selection can sometimes preserve utility at a comparable level, it falls short in achieving effective unlearning. For instance, on Qwen (WMDP), random selection yields a higher utility score (0.5947 vs. 0.5351 for SEUF), but its forget efficacy (FE) remains significantly higher (0.3505 vs. 0.2536 for SEUF), indicating incomplete unlearning. This suggests that selecting the top-1 expert based on affinity scores is crucial for reducing FE while maintaining utility, making it a superior approach to random selection.
Experts with higher affinity scores play a more significant role in unlearning. To further examine the impact of selecting experts based on their affinity scores, we analyze the layer-wise Top-1 expert in DeepSeek on RWKU dataset. In Tab. 5, we present their affinity scores along with the utility (UT) when the expert is involved in unlearning. Due to space constraints, we highlight the top-ranked layer-wise experts (1st to 3rd) and also include several lower-ranked ones (13th to 26th) for comparison. From the results, we observe that the first-ranked expert (with the highest affinity score 0.211) yields the highest UT (0.5485). Overall, UT remains stable at 0.5445 or higher when selecting experts with affinity scores above 0.1. However, when affinity scores drop further (e.g., the 23rd and 26th ranked experts), utility declines more sharply to 0.4262 and 0.2355. These findings emphasize the importance of selecting experts with sufficiently high affinity scores to maintain utility while achieving effective unlearning.
@@
Table 5: Model utility (UT) comparison across unlearned experts with different affinity scores ($s_{i}$) in SEUF+RMU on the RWKU benchmark. UT is compared at a consistent level of forget efficacy (FE $\approx 0.25$).
Rank #1 #2 #3 #13 #20 #23 #26
$s_{i}$ 0.2110 0.1957 0.1695 0.1115 0.0942 0.0844 0.0618
@@
Unlearning resilient to jailbreak attacks. The unlearned model is expected to refuse harmful queries. The forgotten knowledge should not be recovered even through adversarial means. We thus examine the behavior of MoE LLMs unlearned by SEUF under adversarial prompting. Specifically, we test whether SEUF effectively mitigates unauthorized responses by employing the Greedy Coordinate Gradient (GCG) attack Zou et al. (2023) in a white-box setting. This attack optimizes attack prompts to elicit responses that begin with “Sure, here is the answer:”. To increase attack strength, we extend the number of optimization steps to 5,000, while keeping other hyperparameters at their default settings. Given the computational cost ($\sim$ 1 GPU hour on an A100 per soft prompt), we optimize 400 prompts across 400 samples in RWKU for attacking DeepSeek unlearned by SEUF+GA. Since not all responses explicitly begin with "Sure, here is the answer:", we filter for outputs containing the word "answer" and evaluate forget efficacy (FE) both with and without GCG-generated prompts. Our results show that despite being one of the strongest prompt-level attacks, GCG fails to recover forgotten knowledge, as FE remains at 0.01 before and after the attack. To further understand how the GCG attack affects expert selection, we visualize the affinity score of experts in DeepSeek, and compare it with GCG-attacked DeepSeek. Fig. 4 shows that while the GCG attack reduces the affinity score of the target expert, the expert remains ranked as the top-1 in affinity score. This suggests that SEUF maintains stable expert selection even under adversarial influence, ensuring robustness in the unlearning process.
Figure 4: Comparison of affinity scores for all experts in the target layer of DeepSeek unlearned by SEUF + GA on the RWKU dataset, with and without the GCG attack. The target expert is marked as red.
Additionally, we also perform a sensitivity analysis on hyperparameter $\alpha$ in Sec. B in Appendix. The results in Tab. 6 in Appendix indicate that $\alpha=1$ achieves the best performance.
@@ 6 Conclusion
In this paper, we for the first time examine the challenges of applying existing MU techniques to MoE LLMs and carefully investigate the synergy between the dynamic routing system of MoE LLM and the unlearning effects. To address these issues, we proposed SEUF, a novel framework that unlearns most related experts while stabilizing expert selection through a router anchor loss. This approach mitigates expert selection shifts and achieves efficient unlearning with minimal parameter updates. Extensive experiments show that SEUF significantly outperforms traditional unlearning methods and other parameter-efficient fine-tuning techniques, providing a robust solution for MoE LLM unlearning tasks.
7 Limitation
While this study offers valuable insights into unlearning of MoE LLMs, it has certain limitations. First, the evaluation was limited to two datasets due to the scarcity of standardized benchmarks in unlearning. We have used two widely accepted LLM unlearning benchmarks: WMDP (Li et al., 2024) and RWKU (Jin et al., 2024). WMDP. We acknowledge the existence of other commonly used benchmarks, such as TOFU (Maini et al., 2024) and MUSE (Shi et al., 2024). However, these benchmarks are less suitable for our study, as they require models to undergo fine-tuning before unlearning. This additional training step introduces biases in MoE LLMs due to known instability in training, sensitive hyperparameter tuning, and the risk of training collapse (Jiang et al., 2024; Zoph et al., 2022a). These factors make it challenging to isolate the effects of unlearning from the broader impact of model fine-tuning. Expanding the evaluation to a broader range of datasets could enhance the generalizability of the findings. In future work, we plan to explore additional benchmarks, including those that do not require fine-tuning before unlearning, to ensure a more comprehensive assessment of unlearning effectiveness across diverse tasks and model architectures. Second, the study did not apply the unlearning algorithm to Mixtral 8×7B with all parameters unlearned and excluded larger MoE LLM models like DeepSeek-R1 due to computational constraints. Due to the computation limitation, Mixtral is only applied on SEUF and other parameter-efficient fine-tuning unlearning baselines. In future work, we could explore scaling the approach to larger models to evaluate its effectiveness in more complex architectures.
8 Acknowledgement
This work is supported by the NSF award # 2321054.
Appendix A Experiment Setups
Unlearning tasks and datasets. To demonstrate the effectiveness of our proposed method, we evaluate and compare it against different baselines on two widely accepted LLM unlearning benchmarks: WMDP (Li et al., 2024) and RWKU (Jin et al., 2024). WMDP assesses the model’s ability to unlearn and prevent the generation of hazardous knowledge in biosecurity, cybersecurity, and chemical security contexts. RWKU, on the other hand, evaluates the model’s capability to eliminate knowledge about 200 real-world celebrities, simulating a private information protection task. We follow the original study by selecting 100 individuals as unlearning targets. The train_original_passage set, which includes Wikipedia descriptions of these 100 individuals as provided in the paper, is used as the forget set. We note that other commonly used benchmarks, such as TOFU (Maini et al., 2024) and MUSE (Shi et al., 2024), are less appropriate in this work. These benchmarks require models to be fine-tuned before unlearning, which introduces additional biases to the results for MoE LLMs due to the known instability in training and the tricky hyper-parameter tuning involved (Jiang et al., 2024), often leading to training collapse (Zoph et al., 2022a).
Target MoE models to unlearn. We evaluate different unlearning methods on two MoE LLMs: Qwen1.5-MoE-A2.7B-Chat (Qwen), mistralai/Mixtral-8x7B-Instruct-v0.1 (Mixtral), and DeepSeek-V2-Lite (DeepSeek), representing the two mainstream MoE LLM training schemes: upcycle-from-dense and train-from-scratch, respectively. Qwen has a total of 14.3 billion parameters, with 2.7 billion activated during inference, while DeepSeek has 16 billion parameters, of which 2.4 billion are activated during inference. Mixtral has 45 billion parameters, of which 12.9 billion are activated.
Evaluation setup. We evaluate the performance of the unlearned LLMs based on two key metrics: forget efficacy (FE) and preserved model utility (UT). For the WMDP task, FE is measured using the WMDP-Cyber subsets provided by the benchmark. Specifically, we use the accuracy of the forget set after unlearning as the measure of FE. A lower FE indicates better unlearning. Given the four-option multiple-choice format of the test set, the ideal FE is $0.25$, equivalent to random guessing. UT is assessed using the zero-shot accuracy on the MMLU dataset (Hendrycks et al., 2020), which reflects the model’s ability to retain general knowledge. For the RWKU task, we use the Rouge-L recall score to evaluate performance on fill-in-the-blank and question-answer tasks, with lower scores indicating more effective unlearning. Since the task follows a question-answer format, the ideal FE is $0.0$, indicating no overlap between the generated answer and the ground truth. The UT evaluation for RWKU is the same as for WMDP, using the MMLU benchmark. By default, during the unlearning process, we select the model checkpoint that achieves the best balance between FE and UT as the optimal checkpoint.
@@ Appendix A Experiment Setups
We utilize the LM Evaluation Harness (Gao et al., 2024) to measure zero-shot accuracy on the MMLU and WMDP cyber datasets. The mean accuracy across all tasks in MMLU serves as a measure of model utility. For the RWKU dataset, we adhere to the original settings, using the prompt “Please complete the blank in the following question. Question:" for fill-in-the-blank tasks and “Please briefly answer the following question. Question:" for generation tasks.
Unlearning Baselines. We demonstrate the effectiveness of our proposed SEUF framework by comparing it against the LLM unlearning baselines: Gradient Ascent (GA(Eldan and Russinovich, 2023), Gradient Difference (GDiff(Maini et al., 2024) and most recent unlearning algorithm Negative Preference Optimization (NPO(Zhang et al., 2024) and Representation Misdirection for Unlearning (RMU(Li et al., 2024). For each method, we compare the original results with those obtained when incorporating SEUF. Given the parameter efficiency of SEUF, we also compare it with two state-of-the-art parameter-efficient fine-tuning (PEFT) methods for MoE LLMs: the low-rank adaptation scheme (LoRA) (Hu et al., 2021) and the Expert-Specialized Fine-Tuning method (ESFT) Wang et al. (2024b), which is specifically designed for MoE LLMs.
Hyperparameter selection. We consider typical unlearning algorithm as baselines. For RMU, due to the original parameters settings for MoE models fail to unlearn both in DeepSeek and Qwen. We adapt its settings to target all expert MLP layers in fifth, sixth, seventh layers, which align with the settings in the dense model. For the hyperparameters, the retain effect parameter is set to 1200, and $c$ is set to 30000 and 3000 in DeepSeek and Qwen, respectively. We set the learning rate to 5e-5 for GA, NPO, and GD while setting it to 1e-4 for SEUF. The batch size is 4 for GA, NPO, and GD, while it is set to 16 for SEUF. In NPO, the beta value is set to 0.001. The $\lambda$ for the retain loss is set to 1 in both GD and NPO. For RMU, we follow the hyperparameters specified in the original work. We configure the steering coefficients as 8000 for Qwen and 32000 for Deepseek, as SEUF targets deeper layers in these models. For ESFT, we set the threshold $p=0.15$. According to Insight 4 in Sec. 4, we set $M=1$ in the experiment section by default. For LoRA, we apply low-rank adaptation to all layers of the model to enable full-layer fine-tuning. All experiments were conducted in a single run without multiple trials.
Appendix B Sensitivity Analysis of $\alpha$
@@
The hyperparameter $\alpha$ is used for ancho loss in our loss function $\min_{\bm{\theta}}\ell_{f}(\bm{\theta};\mathcal{D}_{f})+\lambda\ell_{r}(\bm{\theta};\mathcal{D}_{r})+\alpha L_{\text{anchor}}^{(l)},$ for introducing the anchor loss $L_{\text{anchor}}$. We conduct experiments on Deepseek unlearned by GA with RWKU dataset to explore the performance of different $\alpha$. As shown in Tab. 6, the results indicate that SEUF is robust to a wide range of $\alpha$ and achieves the best performance when $\alpha=1$.
@@
Appendix C Selection of top $M$ experts in different tasks
We also conduct experiments on Qwen unlearned by GA with RWKU dataset to investigate the optimal selection of $M$. The results in Tab. 7 indicate that SEUF achieves the best performance when only one expert is unlearned $M=1$, which is consistent with the Insight 4.
[…]
Selected experts Top-1 Top-3 Top-6 Top-1+4-shared
Same layer 0.5709 0.3695 0.2572 0.2445
Different layers 0.5709 0.4224 0.3872 -
Appendix D Robustness of Expert Selection
To evaluate the robustness of expert selection under token sampling, we conducted an additional experiment on a consistency analysis on the DeepSeek-V2-Lite model using the WMDP forget set. Specifically, we computed the overlap ratio of selected experts across different token subsets, where overlap is defined as the proportion of shared top-6 experts at each MoE layer.
As shown in Table 8, a subset of 100,000 tokens yields a high overlap (0.94) with the expert selections derived from the full dataset. Furthermore, two independently sampled subsets also show strong agreement with each other (0.87 overlap), indicating that the attribution process is stable across different sampling runs.
Table 8: Expert selection overlap between different sampling splits
Dataset Full Subset 1 Subset 2
Full 1.00 0.94 0.85
Subset 1 0.94 1.00 0.87
Subset 2 0.85 0.87 1.00
Appendix E Experiments on Larger MoE Models
Table 9: Performance of Mixtral 8x7B unlearned by GA on WMDP and RWKU datasets.
Method WMDP RWKU
FE ↓ UT ↑ FE ↓ UT ↑
Pretrained 0.5229 0.6885 0.5820 0.6885
LoRA 0.2658 0.2597 0.0000 0.2295
ESFT 0.2574 0.6386 0.0542 0.6743
SEUF 0.2608 0.6364 0.0455 0.6713
Table 10: Tunable parameter ratio of different methods
Method Tunable Parameter Ratio $\downarrow$
LoRA 0.26%
ESFT 14%
SEUF 0.41%
To explore if SEUF can be applied to larger MoE models, we evaluated SEUF on mistralai/Mixtral-8x7B-Instruct-v0.1 (Mixtral 8x7B), one of the most widely used large-scale open-source MoE models, and compared its performance to other parameter-efficient unlearning baselines.
As shown in Table 9, SEUF achieves comparable or even better utility (UT) while maintaining strong forget efficacy (FE). On the WMDP dataset, SEUF achieves a UT of 0.6364, close to ESFT’s 0.6386 and far better than LoRA’s 0.2597. On RWKU, SEUF reaches 0.6713, again comparable to ESFT (0.6743) and significantly ahead of LoRA (0.2295). Importantly, SEUF does so while updating only 0.41% of parameters, as shown in Table 10, substantially fewer than ESFT’s 14%.
Appendix F The Affect of Weighted Expert Norms
Table 11: Table A: The Spearman’s rank correlation between $g_{i,t}$ and $g_{i,t}E(x_{i})$ for All experts and Top 6 experts across all layers in DeepSeek.
Range All experts Top 6
Correlation 1.0 1.0
As experts may have different weight norms, which could in theory impact the total contribution of their outputs, and that $g_{i,t}E(x_{i})$ might reflect this better than $g_{i,t}$ alone. To investigate this, we computed the Spearman’s rank correlation between $g_{i,t}$ and $g_{i,t}E(x_{i})$ across all MoE layers in DeepSeek using the WMDP dataset. As shown in the Tab 11, the average rank correlation is 1.0 across both all experts and the top-6 experts, indicating that the ordering induced by $g_{i,t}$ closely matches that of $g_{i,t}E(x_{i})$. This suggests that gating scores alone already serve as a strong proxy for expert contribution, even without explicitly incorporating the output norms. This result aligns with the design of the MoE architecture, where routing is learned independently per token while expert weights are optimized to produce scale-compatible outputs under the sparse gating mechanism.